# A novel numerical framework for three-dimensional fully resolved simulation of freely falling particles of arbitrary shape

### v1

@@ 3.4 Spheroids
For the investigation of the rotational motion at finite Reynolds number, the next experiments investigates spheroidal particles. The naming adopted is based on the three orthogonal axes $a$, $b$, $c$ of the spheroids, defining shape parameters: $el$ for the elongation $el=b/a$ and $fl$ for the flatness $fl=c/b$. The two investigated cases are of an oblate spheroid (‘Spheroid-1’, $el=1.00$, $fl=0.20$) and a prolate spheroid (‘Spheroid-2’, $el=0.20$, $fl=1.00$), with detailed parameters provided in Table 2. The numerical output is compared to experimental results of settling particle behavior in a settling chamber setup [66].
@@
In this settling chamber experimental setup [17], particles of different shapes with at least one dimension larger than 100 $\mathrm{\mu}$m are tracked three-dimensionally using four high-speed cameras. The vertical distance of free fall in still air is tracked for a vertical distance of $\sim$60 mm. In this experimental setup, it is observed that the heavy ellipsoidal particles undergo transient oscillations to reach their terminal state [17, 66] and that the microplastic fibers exhibit orientation fluctuations depending on their shape and can also have a preferred orientation [16].
@@
Experimental and numerical observations show that spheroidal particles adopt a preferential orientation on their free-fall trajectory, with the side of largest cross-section facing the downward direction. The particle then oscillates around this orientation. To compare the statistics of the free-fall motion, we compared both the average terminal velocity and the highest oscillation frequency, as shown in Table 4. An outstanding match is observed for both quantities, with an error margin within 1% for the terminal velocity and within 2% for the oscillation frequency.
@@
Table 4: Comparison of the numerical and experimental data of non-spherical spheroid particles. The terminal velocity and the frequency of angular oscillations of the particles from the DNS are shown as $v_{\mathrm{T}}^{D}$ and $f^{D}$ respectively. The mean values of the terminal velocity and the frequency of angular oscillations from the experiments are shown as $\langle V_{\mathrm{T}}^{E}\rangle$ and $\langle f^{E}\rangle$, and their standard deviation as $\sigma(v_{\mathrm{T}}^{E})$ and $\sigma(f^{E})$ respectively. All quantities of the velocity and frequency are in units of m/s and Hz respectively.
3.5 General shape – Lapilli
The final validation case considers a specific volcanic particle of the category of Lapilli, which is collected from a field observation of a volcanic eruption. The particle was digitized and converted to a triangularized surface mesh representation, serving both as input for an arbitrary shape in the virtual wind-tunnel simulator, and as a model for 3D-printed particles for reproducible experimental investigations. The numerical parameters of the particle (volume and detailed tensor of inertia) are computed through numerical integration.
The particle has a diameter of the order of 3 cm and reaches a particle Reynolds number above $40,000$ at its terminal velocity. Three orthogonal side views are shown on Figure 6, together with a snapshot of the virtual free-fall simulation. This case is chosen both to illustrate the capability of the numerical solver to represent particles of arbitrary shape and to produce robust results in a turbulent regime.
Figure 6: Depiction of the volcanic particle of the lapilli family used in the numeric simulation. Image a)-c) show three side views and axis of the $({\bm{n}}_{0},{\bm{n}}_{1},{\bm{n}}_{2})$ reference frame that is attached to the particle. Image d) shows a snapshot of the velocity-norm field on a plane that cuts the particle during the virtual free-fall simulation.
@@ 3.5 General shape – Lapilli
This problem is difficult to tackle experimentally, because of the challenges of reproducing free-fall conditions for this Reynolds number in a laboratory setting. A vertical wind-tunnel is inadequate, in part because of turbulence present in the air inlet. Similarly to the observation at lower Reynolds regimes, the selection of a preferential free-fall orientation, as well as oscillatory and spinning motion around this orientation, dictate the statistics, including the average terminal velocity, of the particle. This behavior relies on a realistic representation of ambient conditions aerodynamic conditions.
@@
In this case, the experiment was conducted in exterior conditions involving a free-fall over the order of 30 m, allowing to measure the average terminal velocity, but without access to additional statistics. This emphasizes the need of a numerical simulation tool to complement laboratory experimentation and provide insight into details of the particle motion.
The numerical simulation is conducted with parameters $d=31.80$ mm, $\kappa=959.32$, $\nu=1.87e^{-5}\mathrm{m^{2}/s}$, $el=0.6$, $fl=0.6$, $w^{*}=30$, $h^{*}=13$, $h_{\mathrm{ini}}^{*}=5$, $h_{\mathrm{end}}^{*}=3$, $l_{\mathrm{sponge}}^{*}=3$, and $N=300$. The particle history consists of a succession of states dominated by an oscillation around a preferred orientation or rotation around a preferred axis. For this reason, the time towards convergence of the statistical average for the terminal velocity is very long. Within the interval from 2s to 4.5 s, an average value of $v_{\mathrm{T}}=-23.5m/s$ is obtained, which differs by 8% from the experimental value of $-21.7m/s$.
All in all, this comparison shows a good match between simulation and experiment, while the simulation reveals further insights into possible modes adopted by the particle, and their impact on the obtained terminal velocity.
@@ 4 Conclusion
This paper presents a state-of-the-art numerical framework for the study of intricate interactions between particles and air under free-fall conditions. One of the most important features of this framework is its ability to represent particles regardless of their shape complexity, thus ensuring a comprehensive representation of real-world scenarios. Another salient feature is the framework’s ability to simulate extended free-fall distances, a feat made possible by a dynamic memory allocation scheme that accounts for the inherent multi-scale nature of such phenomena. These attributes are further strengthened by careful validation against both analytical and experimental datasets, which have shown the high precision and reliability of the tool.
Our numerical investigations have also revealed several key insights. The first is the necessity to resort to higher resolutions than the one expected at first sight. Indeed, the data show that although integral metrics such as total drag can converge at more moderate resolutions, the precise distribution of force over a particle surface, required to determine the torque, requires much higher resolutions. Furthermore, the relative contribution of pressure and shear terms to the force may still vary at increasing resolution, although the total force remains largely invariant. A second interesting observation is that the proper representation of the rotational dynamics of the particles receives critical importance. The oscillatory motion of particles around a particular side has a pronounced effect, with the selected side, or its statistical patterns, playing a central role in determining the free fall velocity.
In essence, the framework presented in this article serves as a solid foundation and invaluable tool for the in-depth exploration of particle free-fall physics. First applications of this framework and through investigations of particle free-fall physics are presented in [66]. Regarding future work, the next steps involve diving into simulations with dynamic mesh refinement. Although such an endeavor comes with substantial technical difficulties, it promises to further improve the resolution close to the particles. In addition, it paves the way for multi-particle simulations, providing a more comprehensive understanding of the dynamics at play.

### v2

@@ 3.4 Terminal velocities of spheroids and transient oscillations
For the validation of angular motion of non-spherical particles at finite Reynolds number, the next set of Palabos Turret simulations investigate the freely falling settling behavior of spheroidal particles. The naming adopted is based on the three orthogonal axes $a$, $b$, $c$ of the spheroids, defining shape parameters: $el$ for the elongation $el=b/a$ and $fl$ for the flatness $fl=c/b$. The two investigated cases are of an oblate spheroid (‘Spheroid-1’, $el=1.00$, $fl=0.20$) and a prolate spheroid (‘Spheroid-2’, $el=0.20$, $fl=1.00$), with detailed parameters provided in Table 2. The numerical output is compared to experimental results of settling particle behavior in the Göttingen Turret settling chamber [69].
@@
In this settling chamber experimental setup [16], particles of different shapes with at least one dimension larger than 100 $\mathrm{\mu}$m are tracked three-dimensionally using four high-speed cameras. The vertical distance of free-fall in still air is tracked for a vertical distance of $\sim$60 mm. In this experimental setup, it is observed that the heavy ellipsoidal particles undergo transient oscillations to reach their terminal state [16, 69] and that the microplastic fibers exhibit orientation fluctuations depending on their shape and can also have a preferred terminal orientation [15, 70].
@@
Experimental and numerical observations show that spheroidal particles adopt a preferred orientation on their free-fall trajectory, with the side of largest cross-section facing the direction of gravity. If a particle is not originally aligned in this preferred direction, it oscillates around and gradually approaches the preferred orientation. To validate the numerical statistics of the freely falling behavior of spheroidal particle, we compared both the terminal velocity and the oscillation frequency with the experimental observations, as shown in Table 4. An excellent match is observed for both quantities, with a relative difference within 1% for the terminal velocity and within 2% for the oscillation frequency.
@@
Table 4: Comparison of the numerical and experimental observation of the settling behavior of non-spherical spheroidal particles. The terminal velocity and the frequency of angular oscillations of the particles from the Palabos Turret simulation are shown as $v_{\mathrm{T}}^{D}$ and $f^{D}$ respectively. The mean values of the terminal velocity and the frequency of angular oscillations from the experiments are shown as $\langle V_{\mathrm{T}}^{E}\rangle$ and $\langle f^{E}\rangle$, and their standard deviation as $\sigma(v_{\mathrm{T}}^{E})$ and $\sigma(f^{E})$ respectively. All quantities of the velocity and frequency are in units of m/s and Hz respectively.
3.5 Free fall of a volcanic particle of non-spherical irregular shape
The final validation case considers a specific volcanic particle of the category of lapilli, which is collected from a field campaign near a volcanic eruption. This particle was imaged with laser line scans under a microscope, and the surface contour points were digitized and converted to a triangulated surface mesh. This surface mesh served both as the input of particle geometry in the Palabos Turret simulation, and as a model for 3D-printing of identical particles for reproducible experimental investigations. This case is chosen both to illustrate the capability of this presented Palabos Turret to simulate particles of arbitrary shape and to produce robust results in a turbulent regime. The dimensions of this particle are 39.72 mm, 28.08 mm and 40.08 mm along the three orthogonal directions as shown in Figure 4(a), with a volume of 1.68e4 mm3 which is equivalent to a sphere of diameter $d_{eq}=$ 31.8 mm.
@@ 3.5 Free fall of a volcanic particle of non-spherical irregular shape
This particle is too large to do laboratory experiment on its free-fall, because of the challenges of reproducing free-fall conditions for high Reynolds number in a laboratory setting. A vertical wind-tunnel is not suitable, especially because of the turbulence introduced through the air inlet, which would unsettle the orientation of the particle. Similarly to the falling behavior of non-spherical particles at lower Reynolds numbers, this particle also exhibits oscillatory and spinning motion in its falling orientation. Such behavior dictates the free-fall statistics, including the average terminal velocity of the particle. A realistic representation of ambient aerodynamic condition is therefore essential.
Figure 4: The Palabos Turret simulation of an irregular particle at $Re\sim 4\times 10^{4}$. (a) 3D model of a volcanic particle captured using a 3D-scanner [71] and a reference normal vector represented with an orange arrow. (b) Time evolution of the angle of the reference normal vector $\theta$ the angle between the normal vector and the perpendicular $y$ direction parallel to $g$. (c) Time evolution of the settling velocity of the particle from the Palabos Turret simulation in orange solid line and a box plot of the terminal velocity of the particle $v_{\mathrm{T}}^{E}$ from experiment. (d) Contours of the magnitude of fluid velocity $|v|$ along a vertical plane passing through the particle. Two streamlines, colored also with $|v|$, captures the complex recirculating nature of the wake.
@@
The experiment was conducted in an exterior setting involving a free-fall over a vertical distance in the order of 30 m. Such large vertical distance allowed to measure the average terminal velocity, but without access to additional small scale statistics. This emphasizes the need of a numerical simulation framework to complement the experimental observation and to provide insight into the details of particle motion. The average terminal velocity obtained experimentally for this particle is $\langle v_{\mathrm{T}}^{E}\rangle=$ 21.36 m/s and it is observed that this particle undergoes rotational motion when settling under gravity. The particle Reynolds number $\mathrm{Re}=\langle v_{\mathrm{T}}^{E}\rangle d_{eq}/\nu$ is $\sim$40000, where the kinematic viscosity of air $\nu$ is 1.73e-5 m2/s.
The Palabos Turret simulation is conducted with parameters $\rho_{\mathrm{f}}=$ 1.096 kg/m3, $\rho_{\mathrm{p}}=$ 1127 kg/m3, $\kappa=\rho_{\mathrm{p}}/\rho_{\mathrm{f}}=1028.28$, $w^{*}=12.58$, $h^{*}=37.74$, $h_{\mathrm{ini}}^{*}=25.16$, $h_{\mathrm{end}}^{*}=3.77$, $l_{\mathrm{sponge}}^{*}=3.02$, and $N=450$. The numerical domain has 273.4 million lattices and was simulated with 2048 MPI (message passing interface) tasks. The time history of the free fall of this particle consists of fluctuations in its orientation, which are shown in Figure 4(b) as a variation of the angle $\theta$ of the normal vector to the vertical direction. Such angular motion changes the projection area of the particle in the falling direction, and the settling velocity of the particle therefore changes with time, as shown in Figure 4(c). During the shown interval of 3 seconds, the average value of the settling velocity – the terminal velocity $\langle v_{\mathrm{T}}^{D}\rangle$ is 22.55 m/s, which differs by 5.3% from the experimental mean value of 21.36 m/s. This comparison shows a good match between the Palabos Turret simulation and experiment, while the simulation reveals further insights into the angular orientations adopted by the particle, and their impact on the settling velocity. Figure 4(d) shows the details of airflow around the particle, as it freely falls under the gravity. The Palabos Turret simulation resolves the details of the boundary layer and the small scale vortices in the recirculating wake behind the particle.
@@ 4 Conclusion
This paper presents a state-of-the-art numerical framework, the Palabos Turret, for studying the falling behaviour of particles of arbitrary shape over a wide range of particle Reynolds numbers spanning 5 orders of magnitude – from the Stokes’ to the Newton’s regime. The recursive-regularized LBM, IBM, and a time-staggered predictor-corrector method are employed to resolve the fluid field, fluid-particle coupling, and particle motion, respectively. A dynamic memory allocation scheme that takes into account the inherent multi-scale nature of the problem allows the simulation of any desired fall distance. Here we have verified the excellent performance of the Palabos Turret with both analytical and experimental data sets.
Although the translational motion of the particles can be correctly captured at fairly coarse grid resolutions, we have found that the surface stresses and rotational dynamics can only be truly captured at fine grid resolutions. For instance, integral metrics such as total drag can converge at lower resolutions, but the accurate force distribution across a particle surface (which is needed to determine the torques) requires much higher resolution. Furthermore, the relative contribution of pressure and shear terms to the force can vary with increasing resolution, even if the total force remains fairly constant. Another interesting observation is that the correct representation of the rotational dynamics of the particles is of critical importance, since the instantaneous settling velocity is a function of particle orientation. Accurate angular motion is required for the actual derivation of the statistics of the settling velocity and lateral movement.
Now that we have verified the reliability of the Palabos Turret, it can be used as an invaluable tool to complement experimental data and to overcome the limitations of experiments in which capturing flow features, surface loads and long particle trajectories are a major challenge. The first application of the Palabos Turret and thorough investigations of the free fall dynamics of ellipsoids at relatively low Reynolds numbers of the order of unity are presented in [69], where further excellent agreement between the Palabos Turret and experimental data is unveiled.
Acknowledgments
T. B. was funded by the German Research Foundation (DFG) Walter Benjamin Position (Project No. 463393443). This work was supported by the Max Planck Society and University of Geneva. Scientific activities are carried out in the Max Planck Institute for Dynamics and Self-Organization (MPIDS) and in University of Geneva. Computational resources from the Max Planck Computing and Data Facility and the MPIDS are gratefully acknowledged. We thank Francesco Marson, Orestis Malaspinas and Eberhard Bodenschatz for support.
Author contributions
Conceptualization and research design: T.B., J.L., Y.W. and G.B.; experiments: T.B.; postprocessing of the experimental data: T.B. and G.B.; simulations: T.B., J.L. and Y.W.; postprocessing of the simulation data: T.B., J.L. and Y.W.; formal analysis: T.B., J.L., Y.W. and G.B.; development of simulation software: T.B., J.L. and Y.W.; visualization: T.B., J.L. and Y.W.; validation: T.B., J.L. and Y.W.; interpretation of results: T.B., J.L., Y.W. and G.B.; writing-original draft: T.B., J.L., Y.W. and G.B; writing-review and editing: T.B., J.L., Y.W. and G.B.
Competing interests
There are no competing interests to declare.
Data and materials availability:
Details of the experiments are available in https://doi.org/10.48550/arXiv.2408.11487. Other data related to this paper may be requested from the authors.
@@ Appendix A Regularized LBM: equilibrium and off-equilibrium
Appendix B Equations for sphere motion
The Basset-Boussinesque-Oseen equation [67]:
Here $r$ is the radius of the sphere.
A solution for rotational motion was first considered by Basset (1888) [72] and later improved by Feuillebois & Lasek (1978) [73]. Here, a constant torque $T$ is applied to a sphere initially at rest. The sphere begins to accelerate angularly and ultimately reaches the terminal angular velocity $\omega$, given by $\omega=T\kappa/(6m\nu)$, where $m$ is the mass of the sphere and $\kappa$ is the density ratio between the sphere and the fluid. The time evolution of the angular velocity $\omega$ is described by the relation:
For a sphere with homogeneous mass distribution, the moment of inertia $I$ takes the value $I=2mr^{2}/5$, and the term $\rho_{\mathrm{f}}/I$ in the equation above is given by $\rho_{\mathrm{f}}/I=15/(8\kappa r^{5}\pi)$.